OpenStudy (anonymous):

Let f(x) = x – 6 and g(x) = x2 – x + 4. Find g(f(x)). x^2 – 13x + 46 x^2 – 11x + 34 x^2 – 2x – 2 x^2 – x – 2

OpenStudy (johnweldon1993):

so you look at your "f" function first....plug "x" in wherever you see an x...sounds pointless i know f(x) = x-6 so now....plug that answer (x-6) into your "g" function wherever you see an x g(x) = x2 – x + 4 g(x-6) = (x-6)² - x-6 + 4 = x² - 12x +36 -x - 6 + 4 = x² -13x + 34

OpenStudy (johnweldon1993):

hmmm that's not right :/ lemme check that over

OpenStudy (anonymous):

yeah i'm not sure, i tried doing it the same way

OpenStudy (johnweldon1993):

oh duh..... lol g(x-6) = (x-6)² - x-6 + 4 = x² - 12x +36 -x + 6 + 4 -12x - 1x is -13x......36..+6...+4....is 46 = x² -13x + 46

OpenStudy (johnweldon1993):

just remember to keep the ( ) around the (x - 6) or else you run into the same mistake I did there so g(x) = x2 – x + 4 g(x-6) = (x-6)² - (x-6) + 4 distribute the - sign into those parenthesis g(x-6) = (x-6)² -x + 6 + 4 and foil out that (x-6)² = x² - 12x +36 -x + 6 + 4 = x² -13x + 46 :)

OpenStudy (johnweldon1993):

see where i made the mistake? and how i fixed it?

OpenStudy (anonymous):

Yes! thank you

OpenStudy (johnweldon1993):

anytime :)